OpenStudy (anonymous):

Find the HCF of 2^120-1 and 2^100-1.

OpenStudy (anonymous):

[(2^100)-1]

OpenStudy (anonymous):

how do u find it?

OpenStudy (anonymous):

[(2^20)-1]

OpenStudy (anonymous):

as per the rules x^n/ x^m = x^n-m so 2^120-100= 2^20-1

OpenStudy (anonymous):

u r chinese right!

OpenStudy (anonymous):

ya

OpenStudy (experimentx):

is that -1 at the end??

OpenStudy (anonymous):

i dont understand the notation is 2^120 - 1 = -1 / 2^120 ?

OpenStudy (experimentx):

looks like it's not easy as it looks \( 2^{120} - 1\) and \( 2^{100} - 1\)

OpenStudy (anonymous):

1 / 2^120 ?

OpenStudy (anonymous):

oh - right - no thats not easy

OpenStudy (anonymous):

hmm..i've think thru it and i've a different explaination: 2^(120)-1= (2^(100)-1)(2^(20)+1) Hence, 2^(120)-1/2^(100)-1= (2^(100)-1)(2^(20)+1)/(2^(100)-1) = 2^(20)+1 (remainder) Thus, 2^(100)-1 is the HCF

OpenStudy (experimentx):

this approach might work. a^n - x^n = (a-x)(a^(n-1) + a^(n-2)x + .... + a x^(n-1) + x^n)

OpenStudy (experimentx):

Oo ... how foolish of me.

OpenStudy (anonymous):

(2^(100)-1)(2^(20)+1) = 2^120 - 2^20 + 2^(100) - 1, not 2^(120)-1

OpenStudy (anonymous):

oo ya, just realised...

OpenStudy (anonymous):

r u sure its not (2^120)^-1 = 1 / 2^120? - in which case its easy to solve.

OpenStudy (anonymous):

in that case the HCF is (2^100)^-1

OpenStudy (experimentx):

@Shu_Yun what is your question?? \( ( 2^{100})^{-1} \) or \( (2^{100})-1\)

OpenStudy (anonymous):

would'nt 2^100 - 1 be prime?? - not sure

OpenStudy (anonymous):

@experimentX (2^100)-1

OpenStudy (experimentx):

(2^100)-1 = (2^20)^50-1^50 = (2^20 - 1){some junk terms} (2^120)-1 = (2^20)^60-1^60 = (2^20 - 1){some junk terms} until now the highest common factor is 2^20 - 1 <--- it might even get higher. keep trying

OpenStudy (anonymous):

i just checked 2^120 - 1 and 2^100 - 1 on wolframalpha - neither are prime

OpenStudy (experimentx):

yeah ... there are some junk terms ... a^n - x^n = (a-x)(a^(n-1) + a^(n-2)x + ..stupid junk terms.. + a x^(n-1) + x^n)

OpenStudy (anonymous):

i've cheated!!!! wolfram came up with 2^20 - 1

OpenStudy (experimentx):

huh??? so my guess was right??

OpenStudy (anonymous):

yes

OpenStudy (experimentx):

well, thanks for checking :)

OpenStudy (anonymous):

wolfram solved arithmetically - writing the values as product of their prime factors

OpenStudy (anonymous):

thats ok